ganeshie8 (ganeshie8):

Prove a number of form \(3^n - 1\) can never be a prime (n > 1 is an integer)

ganeshie8 (ganeshie8):

@Oleg3321

ganeshie8 (ganeshie8):

any ideas ?

ganeshie8 (ganeshie8):

here is a nice property every number of form \(3^n-1\) is divisible by \(2\); that is, an even number

OpenStudy (oleg3321):

ohhhhhh, i see now, so simple lol :)

ganeshie8 (ganeshie8):

How do you know every number of that form is divisible by 2 ?

OpenStudy (oleg3321):

Well when you put any number is the degree of 3 the number will be odd but then you subtract 1 and that makes the number even, which makes it non-prime, correct?

ganeshie8 (ganeshie8):

That's really clever! I had never thought of it that way. great job !

OpenStudy (welshfella):

this problem is deceptively simple.

OpenStudy (mathmate):

:)

ganeshie8 (ganeshie8):

Indeed! Using @Oleg3321 's logic, we can say \(a^n−1\) is never prime for all odd numbers \(a\)

OpenStudy (kainui):

Just cause I like it, \[\frac{3^n-1}{3-1} = \sum_{k=0}^{n-1} 3^k\] So you can write: \[3^n-1 = \sum_{k=0}^{n-1}2*3^k\] which means in base 3, this number looks like: \[(3^n-1)_3=\underset{\text{n times}}{\underbrace{22\cdots22}}\] Just like 100-1=99 same thing. Cute observation I wanted to share. Sometimes you can factor numbers this way easily like this: \[222222 = 2*111111=2*1001*111=2*10101*11\]

ganeshie8 (ganeshie8):

Awesomeness! So that basically works for every integer a>2. Just need to look at the number in base 'a' (a^n - 1)_a = aaa... n times